OpenStudy (anonymous):

heeeellppp.!!!!!!!!! il give medal.!!!... suppose a new car that costs $25000 depreciates by 12% each year. in about how many years will the car be worth $15000? use the equation 15,000=(25000)(0.89)^x and round the value of x to the nearest year?

OpenStudy (anonymous):

?

jigglypuff314 (jigglypuff314):

Hello :) I saw you asking for help in the chat, but I'm not very good at these sorts of problems... :/ I'll do what I can for you though :) 15000=(25000)(0.89)^x divide both sides by 25000 so 0.89^x = 0.6 erm then x ln(0.89) = ln(0.6) so \[x =\frac{ \ln(0.6) }{ \ln(0.89) }\]which you would plug into your calculator :)

OpenStudy (anonymous):

would it be 6? or how do you put it in the calculator and thank you so much.!!!

OpenStudy (anonymous):

so 4?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

thank you .!!

jigglypuff314 (jigglypuff314):

you're welcome :)